# Conditional entropy and information of quantum processes

### v1

@@ Abstract
What would be a reasonable definition of the conditional entropy of bipartite quantum processes, and what novel insight would it provide? We develop this notion using four information-theoretic axioms and define the corresponding quantitative formulas. Our definitions of the conditional entropies of channels are based on the generalized state and channel divergences, such as quantum relative entropy, min- and max-relative entropy, etc. We show that the conditional entropy of quantum channels can potentially reveal important features of the channel, such as its underlying causal structure, which cannot be captured by the entropy of quantum channels or the conditional entropy of bipartite states. Specifically, if the von Neumann conditional entropy $S[A|B]_{\mathcal{N}}$ of a quantum channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ is strictly less than $-\log|A|$, then the channel necessarily has causal influence from $A^{\prime}$ to $B$. Furthermore, we show that our definition of conditional entropy establishes the strong subadditivity of the entropy for quantum channels. We study the total amount of correlations possible due to quantum processes by defining the multipartite mutual information of quantum channels.
@@ I Introduction
Quantum information theory may be considered a fateful marriage between quantum theory [1], one of the two most fundamental theories that describe nature, and the information theory [2]. It allows us to inspect physical processes and the theories describing them from an information-theoretic perspective. Generalization of the information-theoretic tools or concepts like (conditional) entropies of joint probability distributions of random variables to quantum (conditional) entropies of the states of composite quantum systems has proven to be pivotal in both fundamental as well as applied quests [3, 4, 5, 6, 7, 8, 9, 10]. Quantum conditional entropies for states are ubiquitous in studying quantum systems and processes. In a limiting case, we can reduce quantum entropic quantities and their relations to those for random variables and classical processes in (classical) information theory (cf. [11, 7]). However, quantum entropic quantities and their relations offer insights unique to quantum systems (states), i.e., not feasible for (classical) random variables. A quintessential example is the existence of quantum states with negative conditional (von Neumann) entropy [12] with operational meaning in the task of quantum state merging [13, 14]. In contrast, conditional (Shannon) entropy is always nonnegative for joint probability distributions. Negative conditional entropy of bipartite quantum states is the signature of entanglement [12]: no separable states [15] can have negative conditional entropy.
@@
Table 1: We present a comparison between forms of dimensional bounds known for Shannon entropy $H(X)$ and conditional entropy $H(X|Y)$, von Neumann entropy $S(A)_{\rho}$ and conditional entropy $S(A|B)_{\rho}$ for quantum states, and von Neumann entropy $S[A]_{\mathcal{N}}$ and conditional entropy $S[A|B]_{\mathcal{N}}$ of point-to-point channels $\mathcal{N}_{A^{\prime}\to A}$ and bipartite quantum channels $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$, respectively. Shannon entropies can be expressed as Kullback-Leibler (KL) divergence, and von Neumann entropies can be expressed as Umegaki relative entropy, a quantum generalization of KL divergence. Dimensional bounds on the von Neumann conditional entropy of a bipartite channel are from this work, and the remaining are facts known in the literature. Upper bounds for all have the same form. Hence, we may focus on lower bounds, which crudely suggests that lower bound for the von Neumann conditional entropy has no similarity. In contrast, lower bound of the von Neumann entropy of a quantum channel is reflected in lower bound of the von Neumann conditional entropy of a quantum state.
@@
Since quantum channels play crucial roles in understanding the dynamics of quantum systems, characterizing them from an information-theoretic perspective is of paramount importance. This motivates us to formulate conditional entropies of bipartite quantum channels. To achieve this, we take an axiomatic approach similar to the formulation of entropies of quantum channels in [20, 19]. In the limiting scenarios, we can obtain conditional entropies of quantum states and quantum entropies of quantum channels [19] (see also [21, 22, 23]111[22] and [19] independently provided definitions of entropies of a quantum channel, see [19] for comparison between their definitions.). As conditional entropies require at least two random variables or systems, for one to be conditioning (role of side memory) and the other to be nonconditioning, conditional entropy for quantum channels is introduced here for bipartite quantum channels. Lingering questions while conceptualizing the notion of conditional entropies of bipartite quantum channels would naturally be (see Table 1):
@@
  • How novel would the conditional entropy of quantum channels be when compared with the conditional entropy of states? Can the conditional entropy of bipartite quantum channels be strictly lower than the least possible values of entropy of quantum channels? If yes, what would that signature property be in quantum channels?
@@
Axiomatic approaches have been employed to define entropies and conditional entropies of probability distributions and quantum states [24, 25, 26, 7, 27, 28]. Axioms form the minimal set of properties that would make entropy functions meaningful, which could be subject to context. Some of the well-known and widely used examples in classical information theory are Shannon entropy, min-entropy, max-entropy, and Rényi entropies, which have quantum counterparts for states, namely, von Neumann entropy, min-entropy, max-entropy, Rényi entropies, and sandwiched Rényi entropies (see Section III.1.1). There are different ways to generalize classical entropic functions to quantum entropic functions. Some of these families of entropy functions find operational meanings in information-processing tasks that enhance their utility and significance.
@@ II Main results
Conditional entropy of states Conditional entropy of channels
@@
$S(A|B)_{\rho}<0$implies state $\rho_{AB}$ is entangled [12]. $S[A|B]_{\mathcal{N}}<-\log|A|$implies channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ is signalling from $A^{\prime}\to B$.
$S(A|B)<0$ for all pure entangled states $\rho_{AB}$ [12, 13]. $S[A|B]_{\mathcal{N}}<-\log|A|$ for all isomteric channels $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ signalling from $A^{\prime}\to B$.
@@ II.1 Axiomatic approach to conditional entropy of quantum processes
Valid entropy functions of channels have been introduced and discussed in [20, 19] (see also [23]). For the above axioms to make sense, we need a well-meaning definition for the reduced channel, which is a marginal channel analogous to the marginal of states. For any bipartite channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$, a channel $\operatorname{tr}_{B}\circ\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ is its reduced channel on $A$. A reduced channel defined like this is a valid, well-defined channel, as the composition of two channels is always a channel. A reduced channel plays the role of marginal fittingly to define the conditional entropy of the quantum channel and discuss the channel-version of subadditivity or strong subadditivity.
@@ II.2 Formulation of conditional entropies of quantum channels
We formulate the conditional entropies of quantum channels in terms of generalized channel divergence, which in turn are derived from the generalized state divergence. A generalized state divergence [30] is a function $\mathbb{D}(\cdot\|\cdot):{\mathrm{S}t}(A^{\prime})\times{\mathrm{P}os}(A^{\prime})\to\mathbb{R}$ between a quantum state $\rho\in\mathrm{St}(A^{\prime})$ and a positive semidefinite operator $\sigma\in{\mathrm{P}os}(A^{\prime})$ that is monotone (nonincreasing) under the action of quantum channels $\mathcal{N}_{A^{\prime}\to A}$. Based on the state divergence, a divergence function between a channel $\mathcal{N}_{A^{\prime}\to A}$ and a completely positive map $\mathcal{M}_{A^{\prime}\to A}$ is be defined as [31, 23] $\mathbb{D}[\mathcal{N}\|\mathcal{M}]:=\sup_{\rho_{RA^{\prime}}}\mathbb{D}(\mathcal{N}(\rho_{RA^{\prime}})\|\mathcal{M}(\rho_{RA^{\prime}}))$, for states $\rho_{RA^{\prime}}\in{\mathrm{S}t}(RA^{\prime})$. It is convenient for our purpose to express the conditional entropies of quantum channels in terms of divergence functions between a channel and a completely positive map that is monotone under the action of quantum superchannels, i.e., pre-processing and post-processing with quantum channels. We make an observation that a divergence function between a channel and a completely positive map as defined earlier, is always monotone (nonincreasing) under the action of quantum superchannels. In this sense, these divergence functions can be deemed as generalized channel divergence.
Let $\mathcal{R}_{A^{\prime}\to A}$ denote a completely depolarizing linear map that has replacer action, $\mathcal{R}(X_{A^{\prime}})=\operatorname{tr}(X_{A^{\prime}})\mathbbm{1}_{A}$ for all trace-class $X_{A^{\prime}}$, and $\mathbbm{1}_{A}$ is the identity operator. The generalized entropy of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as [19, 23] $\mathbf{S}[A]_{\mathcal{N}}=-\mathbf{D}[\mathcal{N}_{A^{\prime}\to A}\|\mathcal{R}_{A^{\prime}\to A}]$, where $\mathbf{D}[\cdot\|\cdot]$ denotes a generalized channel divergence such that $\mathbf{S}[A]_{\mathcal{N}}$ is a valid entropy function of a channel, as discussed in [20, 19].
@@
$\mathbf{S}[A|B]_{\mathcal{N}}=-\inf_{\mathcal{M}\in{Q}(A^{\prime},A)}\mathbf{D}[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\mathcal{R}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}],$ (1)
@@
$\mathbf{I}[A;B]_{\mathcal{N}}=-\inf_{\mathcal{M}^{1}\in{Q}(A^{\prime},A),\mathcal{M}^{2}\in{Q}(B^{\prime},B)}\mathbf{D}[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\mathcal{M}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}].$ (2)
@@
Interestingly, when the associated generalized divergence is taken to be relative entropy [32], we have ${S}[A|B]_{\mathcal{T}}-{S}[A|BC]_{\mathcal{T}}={I}[A;BC]_{\mathcal{T}}-{I}[A;B]_{\mathcal{T}}$ for tele-covariant tripartite channels [9, Definition 5] $\mathcal{T}_{A^{\prime}BC^{\prime}\to ABC}$.
II.3 von Neumann conditional entropy of bipartite quantum channel
By considering the generalized channel (or state) divergence in Eq. (1) to be Umegaki relative entropy, $D(\rho_{A}\|\sigma_{A}):=\lim_{\varepsilon\to 0^{+}}\operatorname{tr}(\rho_{A}(\log\rho_{A}-\log(\sigma_{B}+\varepsilon\mathbbm{1}_{B})))$, we arrive at the von Neumann conditional entropy $S[A|B]_{\mathcal{N}}$ of a quantum channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$. In general,
@@
For tele-covariant bipartite channels $\mathcal{T}_{A^{\prime}B^{\prime}\to AB}$ [37, 38, 39], i.e., bipartite channels simulable via teleportation with the resource states being their Choi states $\Phi^{\mathcal{T}}_{R_{A}ABR_{B}}:=\mathcal{T}(\Phi_{R_{A}A^{\prime}:R_{B}B^{\prime}})$, where $\Phi_{A:B}$ denotes a maximally entangled state of Schmidt rank $\min\{|A^{\prime}|,|B^{\prime}|\}$ and $R_{A}\simeq A^{\prime},R_{B}\simeq B^{\prime}$,
@@
For a tele-covariant bipartite unitary channel $\mathcal{U}_{A^{\prime}B^{\prime}\to AB}$, we have $S[A|B]_{\mathcal{U}}+\log|A^{\prime}|=S(R_{A}A)_{\Phi^{\mathcal{U}}}=S(R_{B}B)_{\Phi^{\mathcal{U}}}$.

### v2

@@ Abstract
What would be a reasonable definition of the conditional entropy of bipartite quantum processes, and what novel insight would it provide? We develop this notion using four information-theoretic axioms and define the corresponding quantitative formulas. Our definitions of the conditional entropies of channels are based on the generalized state and channel divergences, for instance, quantum relative entropy. We find that the conditional entropy of quantum channels has potential to reveal insights for quantum processes that aren’t already captured by the existing entropic functions, entropy or conditional entropy, of the states and channels. The von Neumann conditional entropy $S[A|B]_{\mathcal{N}}$ of the channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ is based on the quantum relative entropy, with system pairs $A^{\prime},A$ and $B^{\prime},B$ being nonconditioning and conditioning systems, respectively. We identify a connection between the underlying causal structure of a bipartite channel and its conditional entropy. In particular, we provide a necessary and sufficient condition for a bipartite quantum channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ in terms of its von Neumann conditional entropy $S[A|B]_{\mathcal{N}}$, to have no causal influence from $A^{\prime}$ to $B$. As a consequence, if $S[A|B]_{\mathcal{N}}<-\log|A|$ then the channel necessarily has causal influence (signaling) from $A^{\prime}$ to $B$. Our definition of the conditional entropy establishes the strong subadditivity of the entropy for quantum channels. We also study the total amount of correlations possible due to quantum processes by defining the multipartite mutual information of quantum channels.
@@ I Introduction
Quantum information theory may be considered a fateful marriage between quantum theory [1], one of the two most fundamental theories that describe nature, and the information theory [2]. It allows us to inspect physical processes and the theories describing them from an information-theoretic perspective. Generalization of the information-theoretic tools or concepts like (conditional) entropies of joint probability distributions of random variables to quantum (conditional) entropies of the states of composite quantum systems has proven to be pivotal in both fundamental as well as applied quests [3, 4, 5, 6, 7, 8, 9, 10, 11]. The entropy of a system quantifies the amount of uncertainty intrinsic to the system and the conditional entropy quantifies the amount of uncertainty present in the system when partial or side information is provided. Quantum conditional entropies for states are ubiquitous in studying quantum systems and processes. In a limiting case, we can reduce quantum entropic quantities and their relations to those for random variables and classical processes in (classical) information theory (cf. [12, 8]). However, quantum entropic quantities and their relations offer insights unique to quantum systems (states), i.e., not feasible for (classical) random variables. A quintessential example is the existence of quantum states with negative conditional (von Neumann) entropy [13] with operational meaning in the task of quantum state merging [14, 15]. In contrast, conditional (Shannon) entropy is always nonnegative for joint probability distributions. Negative conditional entropy of bipartite quantum states is a signature of entanglement [13]: no separable states [16] can have negative conditional entropy.
@@
Table 1: We present a comparison between the dimensional bounds known for Shannon entropy $H(X)$ and conditional entropy $H(X|Y)$, von Neumann entropy $S(A)_{\rho}$ and conditional entropy $S(A|B)_{\rho}$ for quantum states, and von Neumann entropy $S[A]_{\mathcal{N}}$ and conditional entropy $S[A|B]_{\mathcal{N}}$ of point-to-point channels $\mathcal{N}_{A^{\prime}\to A}$ and bipartite quantum channels $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$, respectively. Shannon entropies can be expressed as Kullback-Leibler (KL) divergence, and von Neumann entropies can be expressed as (Umegaki) relative entropy, a quantum generalization of the KL divergence. Dimensional bounds on the von Neumann conditional entropy of a bipartite channel are from this work, and the remaining are facts known in the literature. Upper bounds for all have the same form. Hence, we may focus on lower bounds for comparison. The lower bound of the von Neumann entropy of a quantum channel is reflected in the lower bound of the von Neumann conditional entropy of a quantum state. In contrast, the lower bound for the von Neumann conditional entropy of a channel stands out from the rest.
@@
Investigating quantum channels from information-theoretic perspective is crucial to understand the dynamics of quantum systems and its utility in technological applications. This motivates us to formulate conditional entropies of bipartite quantum channels. To achieve this, we take an axiomatic approach similar to the formulation of entropies of quantum channels in [21, 20]. In the limiting scenarios, we can obtain conditional entropies of quantum states and quantum entropies of quantum channels [20] (see also [22, 23, 24]111[23] and [20] independently provided definitions of the entropies of a quantum channel, see [20] for comparison between their definitions based on axiomatic approach of [21].). As conditional entropies require at least two random variables or systems, for one to be conditioning (role of side memory) and the other to be nonconditioning, conditional entropy for quantum channels is introduced here for bipartite quantum channels. Lingering questions while conceptualizing the notion of conditional entropies of bipartite quantum channels would naturally be (see Table 1):
@@
  • How novel would the conditional entropy of quantum channels be when compared with the conditional entropy of states? Can the conditional entropy of bipartite quantum channels be strictly lower than the least possible values of entropy of quantum channels? If yes, what would that signature property for quantum channels be?
@@
Axiomatic approaches have been employed to define entropies and conditional entropies of probability distributions and quantum states [25, 26, 27, 8, 28, 29]. Axioms form the minimal set of properties that would make entropy functions meaningful, which could be subject to context. Some of the well-known and widely used examples in classical information theory are Shannon entropy, min-entropy, max-entropy, and Rényi entropies, which have quantum counterparts for states, namely, von Neumann entropy, min-entropy, max-entropy, Rényi entropies, and sandwiched Rényi entropies (see Section III.1.1). There are different ways to generalize classical entropic functions to quantum entropic functions. Some of these families of entropy functions find operational meanings in information-processing tasks that enhance their utility and significance, e.g., [26, 30, 10, 31].
@@ II Main results
Conditional entropy of quantum states Conditional entropy of quantum channels
@@
$S(A|B)_{\rho}<0$implies state $\rho_{AB}$ is entangled [13]. $S[A|B]_{\mathcal{N}}<-\log|A|$implies channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ is signaling from $A^{\prime}\to B$.
$S(A|B)<0$ for all pure entangled states $\rho_{AB}$ [13, 14]. $S[A|B]_{\mathcal{N}}<-\log|A|$ for all isometric channels $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ signaling from $A^{\prime}\to B$.
@@ II.1 Axiomatic approach to conditional entropy of quantum processes
Valid entropy functions of channels have been introduced and discussed in [21, 20] (see also [24]). For the above axiom (P2) to make sense, we need a well-meaning definition for the reduced channel, which is a marginal channel analogous to the marginal of states. For any bipartite channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$, a channel $\operatorname{tr}_{B}\circ\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ is its reduced channel on $A$. A reduced channel defined like this is a valid, well-defined channel, as the composition of two channels is always a channel. A reduced channel plays the role of marginal fittingly to define the conditional entropy of the quantum channel and discuss the strong subadditivity of the channel’s entropy.
@@ II.2 Formulation of conditional entropies of quantum channels
We formulate the conditional entropies of quantum channels in terms of generalized channel divergence, which in turn are derived from the generalized state divergence. A generalized state divergence [33] is a function $\mathbb{D}(\cdot\|\cdot):{\mathrm{S}t}(A^{\prime})\times{\mathrm{S}t}(A^{\prime})\to\mathbb{R}$ between quantum states $\rho,\sigma\in\mathrm{St}(A^{\prime})$ that is monotone (nonincreasing) under the action of quantum channels $\mathcal{N}_{A^{\prime}\to A}$. Based on the state divergence, a divergence function between a channel $\mathcal{N}_{A^{\prime}\to A}$ and a completely positive map $\mathcal{M}_{A^{\prime}\to A}$ is be defined as [34, 24] $\mathbb{D}[\mathcal{N}\|\mathcal{M}]:=\sup_{\rho_{RA^{\prime}}}\mathbb{D}(\mathcal{N}(\rho_{RA^{\prime}})\|\mathcal{M}(\rho_{RA^{\prime}}))$, for states $\rho_{RA^{\prime}}\in{\mathrm{S}t}(RA^{\prime})$. We make an observation that a divergence function between channels as defined earlier, is always monotone (nonincreasing) under the action of quantum superchannels, i.e., pre-processing and post-processing with quantum channels. In this sense, these divergence functions between channels are deemed as generalized channel divergence.
Let $\mathcal{R}_{A^{\prime}\to A}$ denote the completely mixing (depolarizing) linear map that has replacer action, $\mathcal{R}(X_{A^{\prime}})=\operatorname{tr}(X_{A^{\prime}})\mathbbm{1}_{A}$ for all trace-class $X_{A^{\prime}}$, and $\mathbbm{1}_{A}$ is the identity operator. The generalized entropy of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as [20, 24] $\mathbf{S}[A]_{\mathcal{N}}=-\mathbf{D}[\mathcal{N}_{A^{\prime}\to A}\|\mathcal{R}_{A^{\prime}\to A}]$, where $\mathbf{D}[\cdot\|\cdot]$ denotes a generalized channel divergence such that $\mathbf{S}[A]_{\mathcal{N}}$ is a valid entropy function of a channel, see for instance [20].
@@
$\mathbf{S}[A|B]_{\mathcal{N}}=-\inf_{\mathcal{M}\in{Q}(B^{\prime},B)}\mathbf{D}[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\mathcal{R}_{A^{\prime}\to A}\otimes\mathcal{M}_{B^{\prime}\to B}],$ (1)
@@
$\mathbf{I}[A;B]_{\mathcal{N}}=-\inf_{\mathcal{M}^{1}\in{Q}(A^{\prime},A),\mathcal{M}^{2}\in{Q}(B^{\prime},B)}\mathbf{D}[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\mathcal{M}^{1}_{A^{\prime}\to A}\otimes\mathcal{M}^{2}_{B^{\prime}\to B}].$ (2)
@@
Interestingly, when the associated generalized divergence is taken to be relative entropy [35], we have ${S}[A|B]_{\mathcal{T}}-{S}[A|BC]_{\mathcal{T}}={I}[A;BC]_{\mathcal{T}}-{I}[A;B]_{\mathcal{T}}$ for tele-covariant tripartite channels [10, Definition 5] $\mathcal{T}_{A^{\prime}BC^{\prime}\to ABC}$.
II.3 von Neumann conditional entropy and signaling in bipartite quantum channel
By considering the generalized channel (or state) divergence in Eq. (1) to be Umegaki relative entropy, $D(\rho_{A}\|\sigma_{A}):=\lim_{\varepsilon\to 0^{+}}\operatorname{tr}(\rho_{A}(\log\rho_{A}-\log(\sigma_{A}+\varepsilon\mathbbm{1}_{A})))$, we arrive at the von Neumann conditional entropy $S[A|B]_{\mathcal{N}}$ of a quantum channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$. In general,
@@
where $S(A|B)_{\omega}:=-D(\rho_{AB}\|\mathbbm{1}_{A}\otimes\rho_{B})$ is the von Neumann conditional entropy of a state $\omega_{AB}$ and $\omega_{B}:=\operatorname{tr}_{A}(\omega_{AB})$ denotes a reduced state of $\omega_{AB}$. The saturation of the r.h.s. holds, i.e.,
$S[A|B]_{\mathcal{N}}=\inf_{\outerproduct{\psi}{\psi}\in{\mathrm{S}t}(RA^{\prime}B^{\prime})}S(A|RB)_{\mathcal{N}(\outerproduct{\psi}{\psi})}$ (6)
if and only if the bipartite channels $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ are semicausal, i.e., no-signaling from $A^{\prime}\to B$ [38, 39]; whereas, for channels $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ with causal influence, signaling from $A^{\prime}\to B$, the strict inequality “$<$" holds,
@@
For tele-covariant bipartite channels $\mathcal{T}_{A^{\prime}B^{\prime}\to AB}$ [40, 41, 42], i.e., bipartite channels simulable via teleportation with the resource states being their Choi states $\Phi^{\mathcal{T}}_{R_{A}ABR_{B}}:=\mathcal{T}(\Phi_{R_{A}R_{B}:A^{\prime}B^{\prime}})$, where $\Phi_{A:B}$ denotes a maximally entangled state of Schmidt rank $\min\{|A|,|B|\}$ and $R_{A}\simeq A^{\prime},R_{B}\simeq B^{\prime}$,
@@
For a tele-covariant bipartite unitary channel $\mathcal{U}_{A^{\prime}B^{\prime}\to AB}$, we have $S[A|B]_{\mathcal{U}}+\log|A^{\prime}|=-S(R_{A}A)_{\Phi^{\mathcal{U}}}=-S(R_{B}B)_{\Phi^{\mathcal{U}}}$.